zarkam21:

Describe two sets of conditions of heat and work that cause an increase in internal energy.

Vocaloid:

this is pretty difficult to conceptualize but you would either have to increase the heat while keeping it isochoric (W=0) or do work on the gas adiabatically (Q=0)

zarkam21:

Write one or two sentences describing how a heat engine works

zarkam21:

A heat engine converts the flow of heat that results from a difference in temperatures into work..?

Vocaloid:

Good, to be more specific the difference in temps is the difference between the hot and cold reservoirs, the difference in temps causes heat to flow and part of it can be harnessed to do work

zarkam21:

A neatly arranged bookshelf is a very orderly system and has low entropy. Why is it safe to assume that this did not happen naturally? Write one or two sentences explaining your answer

Vocaloid:

Hm. I'm not 100% sure what they're looking for but I'd say that since the # of disorganized states is much much higher than the # of organized states, the probability of it happening naturally is very very low

zarkam21:

During the melting process, how does the energy of the molecules of the substance change as a result of the heat it absorbs?

Vocaloid:

melting = solid --> liquid, do you think energy goes up or down

zarkam21:

down

Vocaloid:

hm, be careful with this one, going from solid --> liquid means the molecules are getting more spread out/moving faster so the energy has to go up

zarkam21:

oh and gas would be the most then right from that logic

Vocaloid:

yes

zarkam21:

How does the heat of fusion of water compare to the heats of fusion of other, similar substances? How does this difference relate to intermolecular forces?

zarkam21:

It is higher because of extensive intermolecular hydrogen bonding.

Vocaloid:

awsome

zarkam21:

How much heat is needed to boil 5.0 kg of water at its boiling point? The latent heat of vaporization for water is 2260 kJ/kg. Show your work.

zarkam21:

Same formula right the q=

Vocaloid:

in this case it would just be heat of vaporization * mass

zarkam21:

2260*5=11300

Vocaloid:

good, just stick the joules unit at the end

zarkam21:

When the temperature is expected to drop below freezing, a farmer places large tubs of water in an unheated food storage shed. Write two or three sentences explaining why the farmer uses this technique to protect the produce from freezing. Hint: Remember that liquid water has more potential energy than ice. (

Vocaloid:

well, the water will freeze - is freezing endo or exothermic?

zarkam21:

endo

Vocaloid:

not quite, for endo it would have to increase in energy meaning it would be melting/evaporating so freezing is exothermic which is why the water will release heat upon freezing thus releasing heat inside the shed and warming it up through convection

zarkam21:

A heat engine converts the flow of heat that results from a difference in temperatures into work. Write one or two sentences describing a very efficient heat engine by comparing the temperatures at which it operates.

Vocaloid:

well if efficiency = 1 - Th/Tc would you want the hot temperature (Th) and the cold temperature (Tc) to be very close or very far apart? as a hint think about what would happen mathematically to the value of 1 - Th/Tc in either case

zarkam21:

um close

Vocaloid:

well, if Th is approximately equal to Tc, then Th/Tc is about 1 so 1 - Th/Tc approaches 0 efficiency so the heat source and cold reservoir need to be far apart in temperature to be efficient

zarkam21:

How could you improve the efficiency of a heat engine by increasing one of the temperatures at which the heat engine operates?

Vocaloid:

well to make Th and Tc farther apart would you increase Th or Tc

zarkam21:

tc

zarkam21:

ugh i dont want to guess

Vocaloid:

well if we assume we have to increase something, then we need to increase Th even higher, if we increase Tc while keeping Th constant they get closer together, so increase Th by adding more fuel/gas/whatever

zarkam21:

okay so improve the efficiency increase Th even higher, if we increase Tc while keeping Th constant they get closer together, so increase Th by adding more fuel/gas/whatever

Vocaloid:

yeah just say increase the hot temperature by adding energy/fuel [I'm not sure about the mechanics of it really]

zarkam21:

Calculate the amount of heat needed to warm 24.6 kg of water from a room temperature of 20.0°C to 60.0°C. Show your work.

Vocaloid:

specific heat 4.186 joule/gram °C convert kg of water to g then use: q = mc*deltaT

Vocaloid:

q = heat = (24.6*1000)g * (4.186 J/gC) * (60-20) = ?

zarkam21:

4119024

Vocaloid:

good since they gave us kg we could do kJ instead of joules and say 4120kJ as the result

zarkam21:

A water heater has an efficiency of 61%. If the water heater takes in 7660 kJ of heat to heat up water for a dishwasher, how much total work does it provide? Show your work.

Vocaloid:

actually nvm it's just e = W/Qin solve for W

zarkam21:

7660000J

Vocaloid:

yeah you can just leave it in kJ, plug in 7660 kJ for Qin and 0.61 for e

zarkam21:

4672.6

Vocaloid:

good, units are kJ b/c the original units were also kJ

zarkam21:

How much energy is lost to the surroundings while heating the water for one cycle of the dishwasher? Show your work

zarkam21:

last one promise

Vocaloid:

is there anything else given in this problem

zarkam21:

no i think it is a continuation of the previous question

Vocaloid:

oh right well if it takes in 7660 kJ and only makes 4672.6 kJ of useful work then everything else must be lost as heat so 7660 - 4672.6 = plug in to a calc and that's your answer

zarkam21:

2987.4

zarkam21:

J ?

Vocaloid:

kJ